# Ortho-Radial Drawing in Near-Linear Time

### v2

@@ 2 Preliminaries
The above lemma shows that $\mathsf{direction}(e^{\star},P,e)\operatorname{mod}4$ is invariant under the choice of the reference path $P$, so the direction of each edge in an ortho-radial representation is well defined, even for the case that $(\mathcal{R},e^{\star})$ might not be drawable. Given the reference edge $e^{\star}$, we let $E_{\mathsf{h}}$ denote the set of all horizontal edges in the east direction, and let $E_{\mathsf{v}}$ denote the set of all vertical edges in the north direction.
@@
Intuitively, $\mathsf{direction}(e,P,e^{\prime})$ quantifies the degree of spirality of $e^{\prime}$ with respect to $e$ and $P$. Unfortunately, Lemma 2.2 does not hold if we replace $\mathsf{direction}(e,P,e^{\prime})\operatorname{mod}4$ with $\mathsf{direction}(e,P,e^{\prime})$. A crucial observation made in [BNRW23] is that such a replacement is possible if we add some restrictions about the positions of $e$, $e^{\prime}$, and $P$. See the following lemma.
Lemma 2.3 ([BNRW23]).
@@
Let $C$ be an essential cycle and let $e$ be an edge in $C$. In view of the above, following [BNRW23], we define the edge label $\ell_{C}(e)$ of $e$ with respect to $C$ to be the value of $\mathsf{direction}(e^{\star},P,e)$, for any choice of reference path $P$ in the exterior of $C$. For the special case that $e=e^{\star}$ and $C=\overline{C_{F_{\mathsf{o}}}}$, we let $\ell_{C}(e)=0$. Intuitively, the value $\ell_{C}(e)$ quantifies the degree of spirality of $e$ from $e^{\star}$ if we restrict ourselves to the exterior of $C$. Consider the edge $e=(u_{1},u_{2})$ in the essential cycle $C=(u_{1},u_{2},u_{3},u_{4})$ in Figure 4 as an example. We have $\ell_{C}(e)=\mathsf{direction}(e^{\star},P,e)=1$, since the reference path $P=(v_{1},v_{5},v_{4},u_{1})$ lies in exterior of $C$.
@@
We are now ready to define the notion of strictly monotone cycles used in (R3). We say that an essential cycle $C$ is monotone if all its edge labels $\ell_{C}(e)$ are non-negative or all its edge labels $\ell_{C}(e)$ are non-positive. Let $C$ be an essential cycle that is monotone. If $C$ contains at least one positive edge label, then we say that $C$ is increasing. If $C$ contains at least one negative edge label, then we say that $C$ is decreasing. Decreasing cycles and increasing cycles are collectively called strictly monotone.
Intuitively, an increasing cycle is like a loop of descending stairs, and a decreasing cycle is like a loop of ascending stairs, so they are not drawable. It was proved in [BNRW23] that $(\mathcal{R},e^{\star})$ is drawable if and only if it does not contain a strictly monotone cycle. Recall again that, throughout the paper, unless otherwise stated, we assume that the given ortho-radial representation already satisfies (R1) and (R2).
Lemma 2.4 ([BNRW23]).
@@
The above theorem improves the previous algorithm of [BNRW23] which costs $O(n^{2})$ time. If the output of $\mathcal{A}$ is a strictly monotone cycle, then the cycle certifies the non-existence of a drawing, by Lemma 2.4. We also extend the above theorem to the case where the reference edge is not fixed.
Theorem 2.
@@
The proof of Theorem 1 is in Section 5, and the proof of Theorem 2 is in Section 6.
@@ 3 Ortho-radial drawings via good sequences
Clearly, if $A=(S_{1},S_{2},\ldots,S_{k})$ is good, then $(S_{1},S_{2},\ldots,S_{i})$ is also good for each $1\leq i<k$. In general, a good sequence might not exist for a given $(\mathcal{R},e^{\star})$. In particular, in order to satisfy (S1), it is necessary that the cycle $\overline{C_{F_{\mathsf{o}}}}$ is a horizontal segment. The sequence $A=(S_{1},S_{2},\ldots,S_{6})$ shown in Figure 6 is a good sequence.
@@
If an ortho-radial drawing of $G_{k}$ for a good sequence $A=(S_{1},S_{2},\ldots,S_{k})$ satisfies (D1), then the drawing also satisfies (D2).
Proof.
See Figure 7 for an illustration of the proof. Suppose (D2) is not satisfied, then we can find three indices $a$, $b$, and $c$ such that the clockwise ordering $(e_{c},e_{b},e_{a})$ given by their $\theta$-coordinates is in the opposite direction of their circular ordering $(e_{a},e_{b},e_{c})$ in $\mathcal{N}_{\mathsf{south}}(A)=(e_{1},e_{2},\ldots,e_{s})$.
Let $G_{k}^{\ast}$ be the graph resulting from identifying the south endpoints of all the edges in $\mathcal{N}_{\mathsf{south}}(A)=(e_{1},e_{2},\ldots,e_{s})$ into a vertex $v^{\ast}$. A planar drawing of $G_{k}^{\ast}$ can be found by extending the given ortho-radial drawing of $G_{k}$ by placing $v^{\ast}$ at the origin and drawing all the edges in $\mathcal{N}_{\mathsf{south}}(A)=(e_{1},e_{2},\ldots,e_{s})$ as straight lines. By (D1), the drawing of $G_{k}^{\ast}$ is crossing-free. Assuming that (D2) is not satisfied, we will derive a contradiction by showing that this drawing cannot be crossing-free, so (D2) must be satisfied.
@@
Consider the cycle $P_{a,b}$ in $G_{k}^{\ast}$. Our assumption on the $\theta$-coordinates for $\{e_{a},e_{b},e_{c}\}$ implies that $e_{c}$ lies in the interior of the cycle $P_{a,b}$ in the above drawing of $G_{k}^{\ast}$. Now consider the path $P_{c,a}$, which starts at $e_{c}$ and ends at $\overline{e_{a}}$. Let $v$ be the first vertex of $P_{a,b}-\{v^{\ast}\}$ that $P_{c,a}$ visits. Since $P_{c,a}$ ends at $\overline{e_{a}}$, such a vertex exists. Let $e$ be the edge incident to $v$ from which $P_{c,a}$ enters $v$. Since $C_{k}^{+}$ is a facial cycle of $G_{k}^{+}$, the circular ordering of the incident edges of $v$ in $C_{k}^{+}$ must respect the counter-clockwise ordering given by $\mathcal{E}(v)$, so $e$ must be an edge in the exterior of the cycle $P_{a,b}$. Therefore, there exist an edge of $P_{c,a}$ and an edge of $P_{a,b}$ crossing each other, since otherwise $P_{c,a}$ cannot go from the interior of $P_{a,b}$ to the exterior of $P_{a,b}$. ∎
@@
For the base case of $k=1$, a good drawing of $G_{1}=S_{1}$ can be constructed, as follows. By (S1), $S_{1}=\overline{C_{F_{\mathsf{o}}}}$, which is the outermost essential cycle. Let $S_{1}=(v_{1},v_{2},\ldots,v_{t})$, where $t=|S_{1}|$ is the number of vertices in the cycle $S_{1}$. Then we may draw $G_{1}=S_{1}$ on the unit circle by putting $v_{j}$ on the point $\left(1,-\frac{j}{t}\cdot 2\pi\right)$, for each $1\leq j\leq t$. The minus sign is due to the fact that $S_{1}$ is oriented in the clockwise direction. The construction of the drawing takes $O\left(|S_{1}|\right)$ time as we need to compute these coordinates. Condition (D1) is satisfied because the drawing does not use any point $(r,\theta)$ with $10\leq r<$ 1.
@@
By (S2), $\mathcal{N}_{\mathsf{north}}(S_{k})$ is a subset of $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k-1})$. For each vertical edge $e_{j}=(x_{j},y_{j})\in\mathcal{N}_{\mathsf{north}}(S_{k})$, We assign the coordinates $(r^{\ast},\theta_{j})$ to $x_{j}$. By (D1), the given drawing does not use any point in $\{(r,\theta)\mid 0\leq r<r_{j}\;\text{and}\;\theta=\theta_{j}\}$, so we may draw $e_{j}$ as a straight line connecting $x_{j}$ and $y_{j}$.
By Lemma 3.1, (D2) follows from (D1). By (D2), for the vertices in $S_{k}$ that we have drawn, that is, the set of vertices in $S_{k}$ that have incident edges in $\mathcal{N}_{\mathsf{north}}(S_{k})$, the clockwise ordering of their $\theta$-coordinates respect the ordering of the horizontal segment $S_{k}=(v_{1},v_{2},\ldots,v_{t})$. If $S_{k}$ is a cycle, then $(v_{1},v_{2},\ldots,v_{t})$ is seen as a circular ordering.
@@
For the case $S_{k}$ is a path, we draw the two endpoints $v_{1}$ and $v_{t}$ of $S_{k}=(v_{1},v_{2},\ldots,v_{t})$, as follows. By (S2), in this case, $\mathcal{N}_{\mathsf{north}}(S_{k})$ is a contiguous subsequence of $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k-1})$. Let $j_{1}$ and $j_{2}$ be the indices such that the subsequence starts at $e_{j_{1}}$ and ends at $e_{j_{2}}$. Let $\epsilon=\min_{1\leq j\leq s}(\theta_{j}-\theta_{j-1})/3$. If $v_{1}$ does not have an incident edge in $\mathcal{N}_{\mathsf{north}}(S_{k})$, then we assign the coordinates $\left(r^{\ast},\theta_{j_{1}}+\epsilon\right)$ to $v_{1}$. Similarly, if $v_{t}$ does not have an incident edge in $\mathcal{N}_{\mathsf{north}}(S_{k})$, then we assign the coordinates $\left(r^{\ast},\theta_{j_{2}}-\epsilon\right)$ to $v_{t}$. Our choice of $\epsilon$ ensures that the range $[\theta_{j_{2}}-\epsilon,\theta_{j_{1}}+\epsilon]$ of radians does not overlap with $\theta_{j}$, for any $e_{j}\in\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k-1})\setminus\mathcal{N}_{\mathsf{north}}(S_{k})$.
@@
For the drawing of $G_{k}$ that we construct, we verify that condition (D1) is satisfied. Consider any vertex $v$ in $G_{k}$ that has an incident edge in $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k})$. Suppose that its coordinates in our drawing are $(r_{v},\theta_{v})$. To prove that (D1) is satisfied, we just need to verify that our drawing does not use any point $(r,\theta)$ with $0\leq r<r_{v}$ and $\theta=\theta_{v}$. For the case $v$ is in $S_{k}$, we have $r_{v}=r^{\ast}$, and our choice of $r^{\ast}$ implies that our drawing does not use any point whose $r$-value is smaller than $r^{\ast}$.
@@
A good drawing of $G_{1}$ can be constructed in $O\left(|S_{1}|\right)$ time. Given a good drawing of $G_{i-1}$, a good drawing of $G_{i}$ can be constructed in $O\left(|S_{i}|\right)$ time. Therefore, given good sequence $A=(S_{1},S_{2},\ldots,S_{k})$, a good drawing of $G_{k}$ can be constructed in $O\left(\sum_{i=1}^{k}|S_{i}|\right)$ time. ∎
@@ 4 Constructing a good sequence
To ensure that a non-empty good sequence exists, by (S1), it is required that $S=\overline{C_{F_{\mathsf{o}}}}$ is a horizontal segment in $\mathcal{S}_{\mathsf{h}}$. If this requirement is not met, then we will add virtual edges to $\mathcal{R}$ to satisfy this requirement. Let $S^{\star}\in\mathcal{S}_{\mathsf{h}}$ be the horizontal segment that contains the reference edge $e^{\star}$, and we have $\mathcal{N}_{\mathsf{north}}(S^{\star})=\emptyset$. We add a virtual horizontal edge $e_{f}$ that connects the two endpoints of $S^{\star}$, so $S^{\star}$ together with $e_{f}$ becomes the new contour of the outer face and is a horizontal segment in $\mathcal{S}_{\mathsf{h}}$. See Figure 8 for an illustration.
Observe that the addition of a virtual edge, in general, does not change the value of edge label $\ell_{C}(e)$ of any edge $e$ in any essential cycle $C$ that already exists in the original graph, as long as the addition of the virtual edge does not destroy (R1) or (R2). The reason is that the calculation of $\ell_{C}(e)$ is invariant under the choice of the reference path $P$ in the calculation of $\ell_{C}(e)$, and there is always a reference path $P$ that already exists in the original graph and does not involve any virtual edge.
@@
Let $S\in\mathcal{S}_{\mathsf{h}}$ such that $\mathcal{N}_{\mathsf{north}}(S)=\emptyset$ and $S\neq\overline{C_{F_{\mathsf{o}}}}$. Such a horizontal segment $S$ can never be added to a good sequence as (S2) requires $\mathcal{N}_{\mathsf{north}}(S)$ to be non-empty. To deal with this issue, we consider the following eligibility criterion for adding a virtual vertical edge incident to such a horizontal segment $S$.
@@
We argue that the addition of $e_{f}=(z,w)$ does not destroy (R1) and (R2). The verification of (R1) is straightforward. We verify (R2) for the case $\mathsf{rotation}(\overline{S}\circ\cdots\circ e^{\prime})=2$, as the other case $\mathsf{rotation}(e^{\prime}\circ\cdots\circ\overline{S})=2$ is similar. The addition of $e_{f}$ decomposes $F$ into two new faces $F_{1}$ and $F_{2}$. Let $F_{1}$ be the one whose facial cycle contains $(u,w,z,x)$ as a subpath, and let $F_{2}$ be the one whose facial cycle contains $(y,z,w,v)$ as a subpath.
@@
Next, consider any $S\in\mathcal{S}_{\mathsf{h}}$ with $|\mathcal{N}_{\mathsf{north}}(S)|\geq 2$. For any two vertical edges $e$ and $e^{\prime}$ such that $e^{\prime}$ immediately follows $e$ in the ordering $\mathcal{N}_{\mathsf{north}}(S)$, we maintain an indicator $X_{e,e^{\prime}}\in\{0,1\}$ such that $X_{e,e^{\prime}}=1$ if $e^{\prime}$ also immediately follows $e$ in $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k})$. Initially, all $X_{e,e^{\prime}}$ are set to 00. For each update to $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k})$, we check and update $X_{e,e^{\prime}}$ for all edges $e$ and $e^{\prime}$ that could be affected. For example, if an edge $\tilde{e}$ is removed from $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k})$, we check if the two edges immediately preceding and following $\tilde{e}$ in $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k})$ belong to $\mathcal{N}_{\mathsf{north}}(S)$ for the same horizontal segment $S$. If the answer is yes, then we set the corresponding indicator $X_{e,e^{\prime}}=1$ for $S$. By (S2), $S$ can be inserted to $A$ if and only if the value of each of its indicators is $1$. Therefore, we can decide whether $S$ should join $W$ by checking the sum $X_{S}$ of all its indicators $X_{e,e^{\prime}}$. This sum $X_{S}$ is updated and checked whenever we update the value of an indicator $X_{e,e^{\prime}}$ for $S$.
@@
The rotation from $e_{i}$ to $e_{j}$ is $r_{j}-r_{i}$ if $1\leq i\leq j\leq s$ and is $r_{j}-r_{i}+4$ if $1\leq j<i\leq s$. Let $i$ be the index such that $e^{\prime}=e_{i}$. Our goal is to search for a horizontal segment $S$ such that $\mathsf{rotation}(e^{\prime}\circ\cdots\circ\overline{S})=2$, so we we may limit our search space to $e_{j}$ such that $r_{j}=r_{i}+2$ or $r_{j}=r_{i}-2$, meaning that we only need to check the two buckets $B_{r_{i}-2}$ and $B_{r_{i}+2}$.
@@
To deal with the above issue, we simply construct the data structure of $F^{\prime}$ from scratch, where the time spent is linear in the number of edges in the facial cycle of $F^{\prime}$. The total cost for the reconstruction throughout the algorithm is upper bounded by $O(s)=O(n)$, where $s$ is the number of edges in $C_{F_{\mathsf{c}}}$ in the original graph $G$. ∎

### v3

@@ 2 Preliminaries
The above lemma shows that $\mathsf{direction}(e^{\star},P,e)\operatorname{mod}4$ is invariant under the choice of the reference path $P$, so the direction of each edge in an ortho-radial representation is well-defined, even for the case that $(\mathcal{R},e^{\star})$ might not be drawable. Given the reference edge $e^{\star}$, we let $E_{\mathsf{h}}$ denote the set of all horizontal edges in the east direction, and let $E_{\mathsf{v}}$ denote the set of all vertical edges in the north direction.
@@
Intuitively, $\mathsf{direction}(e,P,e^{\prime})$ quantifies the degree of spirality of $e^{\prime}$ with respect to $e$ and $P$. Unfortunately, Lemma 2.2 does not hold if we replace $\mathsf{direction}(e,P,e^{\prime})\operatorname{mod}4$ with $\mathsf{direction}(e,P,e^{\prime})$. A crucial observation made in [barth2023topology] is that such a replacement is possible if we add some restrictions about the positions of $e$, $e^{\prime}$, and $P$. See the following lemma.
Lemma 2.3 ([barth2023topology]).
@@
Let $C$ be an essential cycle and let $e$ be an edge in $C$. In view of the above, following [barth2023topology], we define the edge label $\ell_{C}(e)$ of $e$ with respect to $C$ to be the value of $\mathsf{direction}(e^{\star},P,e)$, for any choice of reference path $P$ in the exterior of $C$. For the special case that $e=e^{\star}$ and $C=\overline{C_{F_{\mathsf{o}}}}$, we let $\ell_{C}(e)=0$. Intuitively, the value $\ell_{C}(e)$ quantifies the degree of spirality of $e$ from $e^{\star}$ if we restrict ourselves to the exterior of $C$. Consider the edge $e=(u_{1},u_{2})$ in the essential cycle $C=(u_{1},u_{2},u_{3},u_{4})$ in Figure 4 as an example. We have $\ell_{C}(e)=\mathsf{direction}(e^{\star},P,e)=1$, since the reference path $P=(v_{1},v_{5},v_{4},u_{1})$ lies in exterior of $C$.
@@
We are now ready to define the notion of strictly monotone cycles used in 3. We say that an essential cycle $C$ is monotone if all its edge labels $\ell_{C}(e)$ are non-negative or all its edge labels $\ell_{C}(e)$ are non-positive. Let $C$ be an essential cycle that is monotone. If $C$ contains at least one positive edge label, then we say that $C$ is increasing. If $C$ contains at least one negative edge label, then we say that $C$ is decreasing. Decreasing cycles and increasing cycles are collectively called strictly monotone.
Intuitively, an increasing cycle is like a loop of descending stairs, and a decreasing cycle is like a loop of ascending stairs, so they are not drawable. It was proved in [barth2023topology] that $(\mathcal{R},e^{\star})$ is drawable if and only if it does not contain a strictly monotone cycle. Recall again that, throughout the paper, unless otherwise stated, we assume that the given ortho-radial representation already satisfies 1 and 2.
Lemma 2.4 ([barth2023topology]).
@@
The above theorem improves the previous algorithm of [barth2023topology] which costs $O(n^{2})$ time. If the output of $\mathcal{A}$ is a strictly monotone cycle, then the cycle certifies the non-existence of a drawing, by Lemma 2.4. We also extend the above theorem to the case where the reference edge is not fixed.
Theorem 2.6.
@@
The proof of Theorem 2.5 is in Section 5, and the proof of Theorem 2.6 is in Section 6.
@@ 3 Ortho-radial drawings via good sequences
Clearly, if $A=(S_{1},S_{2},\ldots,S_{k})$ is good, then $(S_{1},S_{2},\ldots,S_{i})$ is also good for each $1\leq i<k$. In general, a good sequence might not exist for a given $(\mathcal{R},e^{\star})$. In particular, in order to satisfy 1, it is necessary that the cycle $\overline{C_{F_{\mathsf{o}}}}$ is a horizontal segment. The sequence $A=(S_{1},S_{2},\ldots,S_{6})$ shown in Figure 6 is a good sequence.
@@
If an ortho-radial drawing of $G_{k}$ for a good sequence $A=(S_{1},S_{2},\ldots,S_{k})$ satisfies 1, then the drawing also satisfies 2.
Proof 3.2.
See Figure 7 for an illustration of the proof. Suppose 2 is not satisfied, then we can find three indices $a$, $b$, and $c$ such that the clockwise ordering $(e_{c},e_{b},e_{a})$ given by their $\theta$-coordinates is in the opposite direction of their circular ordering $(e_{a},e_{b},e_{c})$ in $\mathcal{N}_{\mathsf{south}}(A)=(e_{1},e_{2},\ldots,e_{s})$.
Let $G_{k}^{\ast}$ be the graph resulting from identifying the south endpoints of all the edges in $\mathcal{N}_{\mathsf{south}}(A)=(e_{1},e_{2},\ldots,e_{s})$ into a vertex $v^{\ast}$. A planar drawing of $G_{k}^{\ast}$ can be found by extending the given ortho-radial drawing of $G_{k}$ by placing $v^{\ast}$ at the origin and drawing all the edges in $\mathcal{N}_{\mathsf{south}}(A)=(e_{1},e_{2},\ldots,e_{s})$ as straight lines. By 1, the drawing of $G_{k}^{\ast}$ is crossing-free. Assuming that 2 is not satisfied, we will derive a contradiction by showing that this drawing cannot be crossing-free, so 2 must be satisfied.
@@
Consider the cycle $P_{a,b}$ in $G_{k}^{\ast}$. Our assumption on the $\theta$-coordinates for $\{e_{a},e_{b},e_{c}\}$ implies that $e_{c}$ lies in the interior of the cycle $P_{a,b}$ in the above drawing of $G_{k}^{\ast}$. Now consider the path $P_{c,a}$, which starts at $e_{c}$ and ends at $\overline{e_{a}}$. Let $v$ be the first vertex of $P_{a,b}-\{v^{\ast}\}$ that $P_{c,a}$ visits. Since $P_{c,a}$ ends at $\overline{e_{a}}$, such a vertex exists. Let $e$ be the edge incident to $v$ from which $P_{c,a}$ enters $v$. Since $C_{k}^{+}$ is a facial cycle of $G_{k}^{+}$, the circular ordering of the incident edges of $v$ in $C_{k}^{+}$ must respect the counter-clockwise ordering given by $\mathcal{E}(v)$, so $e$ must be an edge in the exterior of the cycle $P_{a,b}$. Therefore, there exist an edge of $P_{c,a}$ and an edge of $P_{a,b}$ crossing each other, since otherwise $P_{c,a}$ cannot go from the interior of $P_{a,b}$ to the exterior of $P_{a,b}$.
@@
For the base case of $k=1$, a good drawing of $G_{1}=S_{1}$ can be constructed, as follows. By 1, $S_{1}=\overline{C_{F_{\mathsf{o}}}}$, which is the outermost essential cycle. Let $S_{1}=(v_{1},v_{2},\ldots,v_{t})$, where $t=|S_{1}|$ is the number of vertices in the cycle $S_{1}$. Then we may draw $G_{1}=S_{1}$ on the unit circle by putting $v_{j}$ on the point $\left(1,-\frac{j}{t}\cdot 2\pi\right)$, for each $1\leq j\leq t$. The minus sign is due to the fact that $S_{1}$ is oriented in the clockwise direction. The construction of the drawing takes $O\left(|S_{1}|\right)$ time as we need to compute these coordinates. Condition 1 is satisfied because the drawing does not use any point $(r,\theta)$ with $10\leq r<$ 1.
@@
By 2, $\mathcal{N}_{\mathsf{north}}(S_{k})$ is a subset of $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k-1})$. For each vertical edge $e_{j}=(x_{j},y_{j})\in\mathcal{N}_{\mathsf{north}}(S_{k})$, We assign the coordinates $(r^{\ast},\theta_{j})$ to $x_{j}$. By 1, the given drawing does not use any point in $\{(r,\theta)\mid 0\leq r<r_{j}\;\text{and}\;\theta=\theta_{j}\}$, so we may draw $e_{j}$ as a straight line connecting $x_{j}$ and $y_{j}$.
By Lemma 3.1, 2 follows from 1. By 2, for the vertices in $S_{k}$ that we have drawn, that is, the set of vertices in $S_{k}$ that have incident edges in $\mathcal{N}_{\mathsf{north}}(S_{k})$, the clockwise ordering of their $\theta$-coordinates respect the ordering of the horizontal segment $S_{k}=(v_{1},v_{2},\ldots,v_{t})$. If $S_{k}$ is a cycle, then $(v_{1},v_{2},\ldots,v_{t})$ is seen as a circular ordering.
@@
For the case $S_{k}$ is a path, we draw the two endpoints $v_{1}$ and $v_{t}$ of $S_{k}=(v_{1},v_{2},\ldots,v_{t})$, as follows. By 2, in this case, $\mathcal{N}_{\mathsf{north}}(S_{k})$ is a contiguous subsequence of $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k-1})$. Let $j_{1}$ and $j_{2}$ be the indices such that the subsequence starts at $e_{j_{1}}$ and ends at $e_{j_{2}}$. Let $\epsilon=\min_{1\leq j\leq s}(\theta_{j}-\theta_{j-1})/3$. If $v_{1}$ does not have an incident edge in $\mathcal{N}_{\mathsf{north}}(S_{k})$, then we assign the coordinates $\left(r^{\ast},\theta_{j_{1}}+\epsilon\right)$ to $v_{1}$. Similarly, if $v_{t}$ does not have an incident edge in $\mathcal{N}_{\mathsf{north}}(S_{k})$, then we assign the coordinates $\left(r^{\ast},\theta_{j_{2}}-\epsilon\right)$ to $v_{t}$. Our choice of $\epsilon$ ensures that the range $[\theta_{j_{2}}-\epsilon,\theta_{j_{1}}+\epsilon]$ of radians does not overlap with $\theta_{j}$, for any $e_{j}\in\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k-1})\setminus\mathcal{N}_{\mathsf{north}}(S_{k})$.
@@
For the drawing of $G_{k}$ that we construct, we verify that condition 1 is satisfied. Consider any vertex $v$ in $G_{k}$ that has an incident edge in $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k})$. Suppose that its coordinates in our drawing are $(r_{v},\theta_{v})$. To prove that 1 is satisfied, we just need to verify that our drawing does not use any point $(r,\theta)$ with $0\leq r<r_{v}$ and $\theta=\theta_{v}$. For the case $v$ is in $S_{k}$, we have $r_{v}=r^{\ast}$, and our choice of $r^{\ast}$ implies that our drawing does not use any point whose $r$-value is smaller than $r^{\ast}$.
@@
A good drawing of $G_{1}$ can be constructed in $O\left(|S_{1}|\right)$ time. Given a good drawing of $G_{i-1}$, a good drawing of $G_{i}$ can be constructed in $O\left(|S_{i}|\right)$ time. Therefore, given good sequence $A=(S_{1},S_{2},\ldots,S_{k})$, a good drawing of $G_{k}$ can be constructed in $O\left(\sum_{i=1}^{k}|S_{i}|\right)$ time.
@@ 4 Constructing a good sequence
To ensure that a non-empty good sequence exists, by 1, it is required that $S=\overline{C_{F_{\mathsf{o}}}}$ is a horizontal segment in $\mathcal{S}_{\mathsf{h}}$. If this requirement is not met, then we will add virtual edges to $\mathcal{R}$ to satisfy this requirement. Let $S^{\star}\in\mathcal{S}_{\mathsf{h}}$ be the horizontal segment that contains the reference edge $e^{\star}$, and we have $\mathcal{N}_{\mathsf{north}}(S^{\star})=\emptyset$. We add a virtual horizontal edge $e_{f}$ that connects the two endpoints of $S^{\star}$, so $S^{\star}$ together with $e_{f}$ becomes the new contour of the outer face and is a horizontal segment in $\mathcal{S}_{\mathsf{h}}$. See Figure 8 for an illustration.
Observe that the addition of a virtual edge, in general, does not change the value of edge label $\ell_{C}(e)$ of any edge $e$ in any essential cycle $C$ that already exists in the original graph, as long as the addition of the virtual edge does not destroy 1 or 2. The reason is that the calculation of $\ell_{C}(e)$ is invariant under the choice of the reference path $P$ in the calculation of $\ell_{C}(e)$, and there is always a reference path $P$ that already exists in the original graph and does not involve any virtual edge.
@@
Let $S\in\mathcal{S}_{\mathsf{h}}$ such that $\mathcal{N}_{\mathsf{north}}(S)=\emptyset$ and $S\neq\overline{C_{F_{\mathsf{o}}}}$. Such a horizontal segment $S$ can never be added to a good sequence as 2 requires $\mathcal{N}_{\mathsf{north}}(S)$ to be non-empty. To deal with this issue, we consider the following eligibility criterion for adding a virtual vertical edge incident to such a horizontal segment $S$.
@@
We argue that the addition of $e_{f}=(z,w)$ does not destroy 1 and 2. The verification of 1 is straightforward. We verify 2 for the case $\mathsf{rotation}(\overline{S}\circ\cdots\circ e^{\prime})=2$, as the other case $\mathsf{rotation}(e^{\prime}\circ\cdots\circ\overline{S})=2$ is similar. The addition of $e_{f}$ decomposes $F$ into two new faces $F_{1}$ and $F_{2}$. Let $F_{1}$ be the one whose facial cycle contains $(u,w,z,x)$ as a subpath, and let $F_{2}$ be the one whose facial cycle contains $(y,z,w,v)$ as a subpath.
@@
Next, consider any $S\in\mathcal{S}_{\mathsf{h}}$ with $|\mathcal{N}_{\mathsf{north}}(S)|\geq 2$. For any two vertical edges $e$ and $e^{\prime}$ such that $e^{\prime}$ immediately follows $e$ in the ordering $\mathcal{N}_{\mathsf{north}}(S)$, we maintain an indicator $X_{e,e^{\prime}}\in\{0,1\}$ such that $X_{e,e^{\prime}}=1$ if $e^{\prime}$ also immediately follows $e$ in $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k})$. Initially, all $X_{e,e^{\prime}}$ are set to 00. For each update to $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k})$, we check and update $X_{e,e^{\prime}}$ for all edges $e$ and $e^{\prime}$ that could be affected. For example, if an edge $\tilde{e}$ is removed from $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k})$, we check if the two edges immediately preceding and following $\tilde{e}$ in $\mathcal{N}_{\mathsf{south}}(S_{1},S_{2},\ldots,S_{k})$ belong to $\mathcal{N}_{\mathsf{north}}(S)$ for the same horizontal segment $S$. If the answer is yes, then we set the corresponding indicator $X_{e,e^{\prime}}=1$ for $S$. By 2, $S$ can be inserted to $A$ if and only if the value of each of its indicators is $1$. Therefore, we can decide whether $S$ should join $W$ by checking the sum $X_{S}$ of all its indicators $X_{e,e^{\prime}}$. This sum $X_{S}$ is updated and checked whenever we update the value of an indicator $X_{e,e^{\prime}}$ for $S$.
@@
The rotation from $e_{i}$ to $e_{j}$ is $r_{j}-r_{i}$ if $1\leq i\leq j\leq s$ and is $r_{j}-r_{i}+4$ if $1\leq j<i\leq s$. Let $i$ be the index such that $e^{\prime}=e_{i}$. Our goal is to search for a horizontal segment $S$ such that $\mathsf{rotation}(e^{\prime}\circ\cdots\circ\overline{S})=2$, so we may limit our search space to $e_{j}$ such that $r_{j}=r_{i}+2$ or $r_{j}=r_{i}-2$, meaning that we only need to check the two buckets $B_{r_{i}-2}$ and $B_{r_{i}+2}$.
@@
To deal with the above issue, we simply construct the data structure of $F^{\prime}$ from scratch, where the time spent is linear in the number of edges in the facial cycle of $F^{\prime}$. The total cost for the reconstruction throughout the algorithm is upper bounded by $O(s)=O(n)$, where $s$ is the number of edges in $C_{F_{\mathsf{c}}}$ in the original graph $G$.